Paul Coddington Northeast Parallel Architectures Center 111 College Place, 3-228 CST Syracuse University

Phone: (315) 443-2163 FAX: (315) 443-1973 Email: gcf@npac.syr.edu

July 3, 1997

The Graduate Committee School of Computer and Information Science Syracuse University

To whom it may concern,

This is to confirm that Alvin Leung took his written qualifying examination on 10 February 1997. His oral qualifying examination was conducted on 3 July 1997 by Dr Paul Coddington and Professors Geoffrey Fox and Ernie Sibert. Both the written and oral examinations were passed satisfactorily.

Yours sincerely,

Paul Coddington